# A Survey of Secure Semantic Communications

### v1

@@ 1.2 Semantic Communication (SemCom)
Researchers are currently reimagining how mobile communication systems transmit and utilize intelligence information to fully leverage the potential of AI in 6G. An increasing number of researchers are focusing on incorporating semantic information into wireless communications [3] to address this challenge. Shannon and Weaver [4] stated that communications could be categorized into syntactic, semantic, and pragmatic levels. The syntactic level focuses on accurately transmitting the symbols without considering the desired meaning at the semantic level or the intents at the pragmatic level. On the one hand, Semantic Communication (SemCom) enables the communication agents to extract essential semantic information from the original data, thereby reducing network transmission pressure and the processing delays for intelligent tasks [5]. On the other hand, SemCom can be considered a form of AI-induced or brain-like communication mechanism, as the communication is based on a “knowledge base” accumulated through continuous learning [6].
According to recent efforts in SemCom, researchers have mainly focused on mathematical theory studies [7, 8, 9], infrastructure frameworks design [10, 11, 12], joint source-channel coding schemes [13, 14, 15, 16, 17], and advanced semantic extraction techniques for different communication modalities [18, 19, 20, 21]. Niu et al. put forward the system framework of semantic information theory based on synonymous mapping for the first time, which revealed the key relationship between semantic information and grammatical information and expanded the scope of classical information theory in SemCom [7]. In addition, Zhang et al. proposed the model division multiple access technology, which can excavate the model information space based on SemCom to realize the purpose of distinguishing users, and can effectively improve the spectrum efficiency for semantic transmission [10]. Furthermore, Wu et al. introduce a vision transformer-based deep joint source and channel coding (DeepJSCC) scheme for MIMO channels, named DeepJSCC-MIMO, which leverages self-attention to optimize feature mapping and power allocation, enhancing robustness against channel estimation errors and improving both distortion and perceptual quality across diverse scenarios [15]. Moreover, Wang et al. introduce the deep video semantic transmission, a novel method for end-to-end video transmission over wireless channels that leverages DeepJSCC to adaptively transmit video semantic features and utilizes a nonlinear transform and conditional coding to optimize bandwidth use across video frames [19].
@@ 1.3 Secure SemCom
From a physical-layer security (PLS) perspective, Li et al. [32] developed a deep neural network-based secure SemCom system called DeepSSC, which uses a two-phase training strategy. Similarly, Mu et al. [33] proposed another PLS approach that leverages semantic streams as beneficial noise to interfere with eavesdroppers, optimizing power allocation and decoding order to enhance secrecy. In the realm of learning-based techniques, Liu et al. [34] introduced a method using the soft actor-critic algorithm to optimize power control in covert SemCom. Qin et al. [35] suggested generating a semantic key and employing subcarrier-level obfuscation to strengthen security in static channels. For image transmission, Tang et al. [36] employed signal steganography by embedding private image signals into non-sensitive host images to ensure that only legitimate receivers can extract private data. Lastly, Xu et al. [37] proposed a covert SemCom framework for wireless edge networks that uses a full-duplex receiver to generate artificial noise, thereby interfering with eavesdroppers.
Expanding on these efforts, a comprehensive survey by Won et al. [38] provides an in-depth exploration of resource management, security, and privacy challenges in SemComs. The survey highlights the unique vulnerabilities of SemCom systems, including encoder or decoder attacks, knowledge base manipulation, and transmission channel security issues. To address these, the authors outline effective countermeasures such as adversarial training, differential privacy, and quantum key distribution. Notably, the survey also emphasizes the importance of semantic-aware metrics and performance optimization techniques to enhance system resilience against eavesdropping, jamming, and semantic noise. This work serves as a foundational reference for integrating robust security and privacy mechanisms into future SemCom systems.
@@
Representative overview/survey papers on secure SemCom are listed in Tab. LABEL:tab1. Won et al. [38] comprehensively review resource management, security, and privacy in SemCom, introducing semantic-aware resource allocation and identifying open research challenges. Guo et al. [24] survey SemCom networks, addressing multi-layered architecture, security, and privacy issues while proposing future directions. E. Sagduyu et al. [39] analyze deep learning (DL)-based vulnerabilities, particularly adversarial attacks, and propose novel models to enhance system resilience. Du et al. [40] revisit wireless security in the Semantic IoT, proposing new metrics like semantic secrecy outage probability and highlighting key differences from traditional IoT. Shen et al. [41] discuss secure SemCom challenges and propose defenses such as zero-knowledge proofs. Yang et al. [42] examine SemCom security fundamentals, emphasizing protection mechanisms for semantic transmission and machine learning (ML) models. Luo et al. [43] review SemCom theories and advancements, focusing on challenges in semantic extraction, compression, and transmission. He et al. [44] introduce a mixture-of-experts model for 6G SemCom, showcasing its success in mitigating heterogeneous attacks and preserving task performance.
@@ 1.4 Motivations
Despite the fact that many researchers concentrate on ensuring the security of SemCom, it is surprising to find that a comprehensive understanding of the state-of-the-art in secure SemCom and its fundamental principles remains elusive. Furthermore, insufficient progress has been made in understanding the security and privacy risks associated with SemCom. Therefore, the primary aim of this paper is to provide a detailed survey of the characteristics and technologies that can be utilized in the field of secure SemCom. We hope that this work will provide researchers with a foundational understanding and inspiration to further advance the development of SemCom.
@@ 1.5 Contributions
    Initially, we provide an overview of SemCom architecture from both link and network levels. The link-level architecture focuses on the joint optimization of each module, while the network-level architecture highlights the transmission process of deep joint source and channel coding (JSCC) models.
@@
    Following this, we outline the entire life cycle of SemComs, encompassing model training, model transfer, and semantic information transmission stages. We follow the measures of security and privacy in AI field, including confidentiality, integrity, and availability, to describe SemCom systems. The security and privacy challenges in these stages include: 1) model training, which faces knowledge base poisoning attacks, gradient leakage, server vulnerabilities, and attacks targeting communication bottlenecks; 2) model transfer, which is susceptible to model slice attacks; and 3) semantic information transmission, which is vulnerable to semantic adversarial attacks, semantic eavesdropping attacks, semantic inference attacks, and semantic jamming attacks.
@@
    To counteract these security and privacy threats in SemCom, we summarize various defense technologies. These include: 1) data cleaning to address dirty data that may introduce noise, missing values, or inconsistencies; 2) robust learning to enhance the resistance of DL models to noise or input disturbances; 3) defensive methods against backdoor attacks, provided from both data and model perspectives; 4) adversarial training, approached from sample generation and model optimization angles; 5) differential privacy, which introduces noise into the data to prevent attackers from accurately restoring individual information; 6) cryptography technology, further classified into homomorphic encryption, secure multi-party computation, trusted execution environments, and secure aggregation; 7) blockchain technology, which combines data blocks into a tamper-proof and unforgeable chain structure; 8) model compression, further categorized into model pruning, parameter quantization, low-rank decomposition, and knowledge distillation; and physical-layer security, which includes beamforming, artificial noise, relay cooperation, intelligent reflecting surfaces, physical-layer key generation, and physical-layer authentication.
@@ 1.6 Organizations
DNN Deep Neural Network PHE Partially Homomorphic Encryption
@@ 2.1 Link-Level Architecture of SemCom
Figure 2: Comparison of the link-level architecture between traditional communication and SemCom systems.
@@
As shown in Fig. 2(b), the information source is directly processed by the semantic encoder in the semantic space, which aims to extract the contextual meaning of the input message. Then the extracted semantic symbols are passed into the physical channel, where the semantic soft matching is achieved to adapt to both the source information and the channel noise. At the receiver side, the noised semantic symbols are further processed by the semantic decoder, which recovers the source information under the joint consideration of channel noise and the received semantics. Since the semantic encoder and the semantic decoder are jointly trained under the semantic space and adapted according to the propagating channel environment, it eliminates the environment’s effect on source information recovery facilitation under the receiver semantic space. It is noted that the multi-modal or heterogeneous original information requires combining other technologies such as semantic synthesis of multiple sources to deliver the desired meaning and match the transmission goal efficiently.
@@ 2.2 Network-Level Architecture of SemCom
Figure 3: Network-level architecture of SemCom systems.
@@
    The functional nodes on the first layer (cloud) consistently generate highly effective models trained on vast databases collected in a distributed manner, ensuring remarkable efficiency, accuracy, robustness, and reliability. For instance, Google’s servers host advanced models like the Switch Transformer, exemplifying state-of-the-art performance in natural language processing [51]. Moreover, large language models (LLM) provided by AI companies including OpenAI’s ChatGPT and Anthropic’s Claude can also be deployed in the cloud to offer intelligent semantic parsing and generation services [52]. Additionally, cloud platforms can utilize incentive mechanisms, such as blockchain-based reward schemes, to encourage user engagement and contributions to these large AI models.
@@
    The nodes with medium calculation and storage capability, such as mobile edge computing (MEC), can be placed on the second layer (edge). Upon receiving the model from the functional nodes in the first layer, the nodes on the second layer can retrain the model via the techniques such as meta-learning [53], transfer learning [54], knowledge distillation [55], model distillation [56] and federated learning (FL) [57] for different purposes. For example, the edge nodes utilize distributed learning techniques to achieve global collaboration across different knowledge base. By leveraging data with heterogeneous semantic distributions from geographically dispersed locations, the collaboration among edge nodes mitigates the impact of biased data samples on AI model performance, thereby constructing a globally consistent semantic knowledge base [57]. Besides, the edge nodes can compress the large AI models from the first layer to varying degrees based on dynamic channel conditions and resource constraints, making it feasible to transmit the AI models to the nodes on the third layer within the limited bandwidth, storage and computation resources.
@@ 3 Security and Privacy Breaches of SemCom
This section first introduces the full life cycle of SemCom, which includes model training, model transfer, and semantic information transmission phases. Then, it elaborates on the potential security and privacy threats that may arise in each of these three stages.

### v2

@@ 1.2 Semantic Communication (SemCom)
Researchers are currently reimagining how mobile communication systems transmit and utilize intelligence information to fully leverage the potential of AI in 6G. An increasing number of researchers are focusing on incorporating semantic information into wireless communications [3] to address this challenge. Shannon and Weaver [4] stated that communications could be categorized into syntactic, semantic, and pragmatic levels. The syntactic level focuses on accurately transmitting the symbols without considering the desired meaning at the semantic level or the intents at the pragmatic level. On the one hand, Semantic Communication (SemCom) enables the communication agents [^1] In SemComs, “communication agents” refer to the intelligent entities or communication entities that participate in the communication process. They leverage advanced AI technologies and specific functional modules to achieve accurate transmission and understanding of semantic information. to extract essential semantic information from the original data, thereby reducing network transmission pressure and the processing delays for intelligent tasks [5]. On the other hand, SemCom can be considered a form of AI-induced or brain-like communication mechanism, as the communication is based on a “knowledge base” accumulated through continuous learning [6].
According to recent efforts in SemCom, researchers have mainly focused on mathematical theory studies [7, 8, 9], infrastructure frameworks design [10, 11, 12], joint source-channel coding schemes [13, 14, 15, 16, 17], and advanced semantic extraction techniques for different communication modalities [18, 19, 20, 21]. Niu et al. put forward the system framework of semantic information theory based on synonymous mapping for the first time, which revealed the key relationship between semantic information and grammatical information and expanded the scope of classical information theory in SemCom [7]. In addition, Zhang et al. proposed the model division multiple access technology, which can excavate the model information space based on SemCom to realize the purpose of distinguishing users, and can effectively improve the spectrum efficiency for semantic transmission [10]. Furthermore, Wu et al. introduce a vision transformer-based deep joint source and channel coding (JSCC) scheme for MIMO channels, named DeepJSCC-MIMO, which leverages self-attention to optimize feature mapping and power allocation, enhancing robustness against channel estimation errors and improving both distortion and perceptual quality across diverse scenarios [15]. Moreover, Wang et al. introduce the deep video semantic transmission, a novel method for end-to-end video transmission over wireless channels that leverages deep JSCC to adaptively transmit video semantic features and utilizes a nonlinear transform and conditional coding to optimize bandwidth use across video frames [19].
@@ 1.3 Secure SemCom
From a physical-layer security (PLS) perspective, Li et al. [32] developed a deep neural network (DNN)-based secure SemCom system called DeepSSC, which uses a two-phase training strategy. Similarly, Mu et al. [33] proposed another PLS approach that leverages semantic streams as beneficial noise to interfere with eavesdroppers, optimizing power allocation and decoding order to enhance secrecy. In the realm of learning-based techniques, Liu et al. [34] introduced a method using the soft actor-critic algorithm to optimize power control in covert SemCom. Qin et al. [35] suggested generating a semantic key and employing subcarrier-level obfuscation to strengthen security in static channels. For image transmission, Tang et al. [36] employed signal steganography by embedding private image signals into non-sensitive host images to ensure that only legitimate receivers can extract private data. Lastly, Xu et al. [37] proposed a covert SemCom framework for wireless edge networks that uses a full-duplex receiver to generate artificial noise, thereby interfering with eavesdroppers.
@@
Representative overview/survey papers on secure SemCom are listed in Tab. LABEL:tab1. Won et al. [38] comprehensively review resource management, security, and privacy in SemCom, introducing semantic-aware resource allocation and identifying open research challenges. This survey also emphasizes the importance of semantic-aware metrics and performance optimization techniques to enhance system resilience against eavesdropping, jamming, and semantic noise [^2] Semantic noise can originate from the following aspects. (1) Semantic ambiguity is a common source of semantic noise in text. Minor changes in words or adjustments to sentence structures can potentially lead to misunderstandings of the transmitter’s intent by the receiver. Additionally, different cultural backgrounds and language habits may result in diverse interpretations of the same information, thereby generating semantic noise. (2) Transmission errors, due to factors such as channel noise, can alter the semantic content of the information. (3) Data compression and encoding may introduce a certain level of semantic noise as they may lead to the loss or distortion of some details in the information. (4) Adversarial attacks refer to situations where malicious attackers may intentionally introduce semantic noise to disrupt or sabotage the communication process.. Guo et al. [24] survey SemCom networks, addressing multi-layered architecture, security, and privacy issues while proposing future directions. E. Sagduyu et al. [39] analyze deep learning (DL)-based vulnerabilities, particularly adversarial attacks, and propose novel models to enhance system resilience. Du et al. [40] revisit wireless security in the Semantic IoT, proposing new metrics like semantic secrecy outage probability and highlighting key differences from traditional IoT. Shen et al. [41] discuss secure SemCom challenges and propose defenses such as zero-knowledge proofs. Yang et al. [42] examine SemCom security fundamentals, emphasizing protection mechanisms for semantic transmission and machine learning (ML) models. Luo et al. [43] review SemCom theories and advancements, focusing on challenges in semantic extraction, compression, and transmission. He et al. [44] introduce a mixture-of-experts model for 6G SemCom, showcasing its success in mitigating heterogeneous attacks and preserving task performance.
@@
/ Our Survey Provide an overview of SemCom architecture from both link and network levels, outline the entire life cycle of SemComs, summarize various defense technologies, and point out the future research directions.
@@ 1.4 Motivations
Despite the fact that many researchers concentrate on ensuring the security of SemCom, it is surprising to find that a comprehensive understanding of the state-of-the-art in secure SemCom and its fundamental principles remains elusive. For example, [24] and [38] do not provide the basic principle or detailed classification of defense technologies. [39] only discusses adversarial attacks but does not provide a comprehensive overview of other attack threats and their corresponding defense methods. Therefore, the main objective of this paper is to offer an in-depth exploration of the characteristics and technologies that can be employed in the realm of secure SemCom. Initially, we introduce the architectures of SemCom at both the link and network levels for readers unfamiliar with the concept. Subsequently, we delineate the life cycle of SemCom and discuss the security and privacy concerns associated with each stage. Furthermore, we present various defensive techniques, some of which have already been implemented in SemCom, while others hold promise for its future, to facilitate a better understanding for readers without a background in security and privacy. Finally, we outline future research directions, highlighting research challenges and potential solutions. We hope that this work will provide researchers with a foundational understanding and inspiration to further advance the development of SemCom.
@@ 1.5 Contributions
    Initially, we provide an overview of SemCom architecture from both link and network levels. The link-level architecture focuses on the joint optimization of each module, while the network-level architecture highlights the transmission process of deep JSCC models.
@@
    Following this, we outline the entire life cycle of SemComs, encompassing model training, model transfer, and semantic information transmission stages. We follow the measures of security and privacy in AI field, including confidentiality, integrity, and availability, to describe SemCom systems. The security and privacy challenges in these stages include: 1) model training, which faces knowledge base poisoning attacks, gradient leakage, server vulnerabilities, and attacks targeting communication bottlenecks; 2) model transfer, which is susceptible to model slice availability attacks and model slice forgery attacks; and 3) semantic information transmission, which is vulnerable to semantic adversarial attacks, semantic eavesdropping attacks, semantic inference attacks, and semantic jamming attacks.
@@
    To counteract these security and privacy threats in SemCom, we summarize various defense technologies. These include: 1) data cleaning to address dirty data that may introduce noise, missing values, or inconsistencies; 2) robust learning to enhance the resistance of DL models to noise or input disturbances; 3) defensive methods against backdoor attacks, provided from both data and model perspectives; 4) adversarial training, approached from sample generation and model optimization angles; 5) differential privacy, which introduces noise into the data to prevent attackers from accurately restoring individual information; 6) cryptography technology, further classified into homomorphic encryption, secure multi-party computation, trusted execution environments, secure aggregation, and quantum cryptography; 7) blockchain technology, which combines data blocks into a tamper-proof and unforgeable chain structure; 8) model compression, further categorized into model pruning, parameter quantization, low-rank decomposition, and knowledge distillation; and physical-layer security, which includes beamforming, artificial noise, relay cooperation, intelligent reflecting surfaces, physical-layer key generation, and physical-layer authentication.
@@ 1.6 Organizations
DNN DNN PHE Partially Homomorphic Encryption
@@ 2.1 Link-Level Architecture of SemCom
Figure 2: Comparison of the link-level architecture between traditional communication and SemCom systems, where (a) shows the architecture of the traditional communication system, where the source and channel encoders operate independently to transmit and recover bits through the physical channel, focusing solely on accurate bit replication without considering semantic content; (b) illustrates the architecture of the SemCom system, where the semantic encoders and decoders are jointly trained to enable efficient transmission and recovery of semantically meaningful information, adapting to both the source and channel conditions. The above procedure is achieved by the transmission of models, which serve as semantic knowledge base for synonymous mapping, recovery and generation.
@@
As shown in Fig. 2(b), the information source is directly processed by the semantic encoder in the semantic space, which aims to extract the contextual meaning of the input message. Then the extracted semantic symbols are passed into the physical channel, where the semantic soft matching 333“Semantic soft matching” refers to a more flexible and tolerant matching approach adopted in the process of SemComs to process the semantic content of information. Compared to strict, rigid matching criteria in traditional communications, soft matching may allow for a certain degree of semantic differences or ambiguity, as long as these differences do not impair the overall understanding and transmission of the information. is achieved to adapt to both the source information and the channel noise. At the receiver side, the noised semantic symbols are further processed by the semantic decoder, which recovers the source information under the joint consideration of channel noise and the received semantics. Since the semantic encoder and the semantic decoder are jointly trained under the semantic space and adapted according to the propagating channel environment, it eliminates the environment’s effect on source information recovery facilitation under the receiver semantic space. It is noted that the multi-modal or heterogeneous original information requires combining other technologies such as semantic synthesis of multiple sources to deliver the desired meaning and match the transmission goal efficiently.
@@ 2.2 Network-Level Architecture of SemCom
Figure 3: Network-level architecture of SemCom systems, where the semantic collaboration spans across three layers: cloud, edge, and device. The first layer (cloud) hosts highly efficient models trained on vast distributed databases, providing intelligent semantic services and incentivizing user contributions. The second layer (edge) includes mobile edge computing nodes that retrain models using techniques like meta-learning and federated learning, collaborating globally to mitigate data bias and adapt to dynamic conditions. The third layer (device) consists of resource-constrained end nodes that retrieve and update models from the edge, ensuring synchronization for semantic recovery during transmission.
@@
    The functional nodes on the first layer (cloud) consistently generate highly effective models trained on vast databases collected in a distributed manner, ensuring remarkable efficiency, accuracy, robustness, and reliability. For instance, Google’s servers host advanced models like the Switch Transformer, exemplifying state-of-the-art performance in natural language processing [51]. Moreover, large language models (LLM) provided by AI companies, including OpenAI’s ChatGPT, Anthropic’s Claude, and the recently popular DeepSeek, can also be deployed in the cloud to offer intelligent semantic parsing and generation services [52]. Additionally, cloud platforms can utilize incentive mechanisms, such as blockchain-based reward schemes, to encourage user engagement and contributions to these large AI models.
@@
    The nodes with medium calculation and storage capability, such as mobile edge computing (MEC), can be placed on the second layer (edge). Upon receiving the model from the functional nodes in the first layer, the nodes on the second layer can retrain the model via the techniques such as meta-learning [53], transfer learning [54], knowledge distillation [55], model distillation [56] and federated learning (FL) [57] for different purposes. For example, the edge nodes utilize distributed learning techniques to achieve global collaboration across different knowledge base. By leveraging data with heterogeneous semantic distributions [^4] Information may originate from diverse sources such as text, images, audio, video, and more. These different information sources may adopt varying methods when expressing semantics, leading to heterogeneity in semantic distribution. Even for the same information source, different individuals or systems may utilize different modes of expression or linguistic styles to convey information. This variability in expression similarly results in heterogeneity in semantic distribution. from geographically dispersed locations, the collaboration among edge nodes mitigates the impact of biased data samples on AI model performance, thereby constructing a globally consistent semantic knowledge base [57]. Besides, the edge nodes can compress the large AI models from the first layer to varying degrees based on dynamic channel conditions and resource constraints, making it feasible to transmit the AI models to the nodes on the third layer within the limited bandwidth, storage and computation resources.
@@ 3 Security and Privacy Breaches of SemCom
This section first introduces the full life cycle of SemCom, which includes model training, model transfer, and semantic information transmission phases. Then, it elaborates on the potential security and privacy threats that may arise in each of these three stages. The detailed security and privacy threats are listed in Tab. 3.
Table 3: List of Security and Privacy Threats in SemCom.
Attacks Positions Phases Properties Effects
Knowledge base poisoning attacks Knowledge base Model training Integrity and Availability It entails the premeditated introduction of harmful data into the knowledge base by an adversary, particularly in a publicly accessible one. The extraction and restoration of semantic characteristics hinge on the quality of the data in the knowledge base, and contaminated data undermines the authenticity and accessibility of the information. It is divided into data poisoning attacks, model poisoning attacks, and backdoor attacks.
Gradient leakage Model transmission Model training Confidentiality The attacker can update the dummy input and output to minimize the gradient distance. The dummy output is the same size as the dummy input during the deep JSCC model training instead of a class probability vector during the classification model training. Despite the difficulty of stealing data in realistic training regimes, the deep JSCC model’s collaborative training procedure may lead to a deeper and easier privacy leakage.
Vulnerability of servers Model transmission Model training Confidentiality, Integrity, and Availability If the server’s security is inadequate, attackers may directly access the global model from the server, further tampering with the model or extracting sensitive information, which poses a serious threat to the integrity and accuracy of the SemCom model. If the server is attacked during the aggregation process, it may lead to tampering with the model updates, thereby affecting the training effectiveness and performance of the global model.
Attacks against communication bottlenecks Model transmission Model training Availability Despite they infrequent occurrence, the consequences they trigger can be severe, potentially causing significant disruptions to the federated training environment of SemCom models. For exmaple, models have become increasingly intricate, with a substantial increase in the number of model parameters that change during each iteration, thereby significantly elevating communication overhead. Furthermore, limited network bandwidth can lead to delays for edge servers when uploading local model updates to the cloude server or downloading the global model, resulting in dropout.
Model slice availability attacks Model transmission Model transfer Availability Model slice availability attacks can corrupt the parameters of the model slices, directly impacting their SemCom performance once deployed.
Model slice forgery attacks Model transmission Model transfer Integrity Model slice forgery attacks involve deploying model parameters pre-trained by the attacker at the terminal, thereby steering the user’s SemCom results towards the direction desired by the attacker.
Semantic adversarial attacks Source of information and physical channel Semantic information transmission Availability SemCom systems on DL models for the extraction, encoding, and transmission of semantic information. Semantic adversarial attacks are covert and efficient Attackers can intentionally introduce imperceptible perturbations into the input data, causing the model to make erroneous predictions or exhibit unintended behaviors. Additionally, by introducing meticulously crafted perturbations into the channels, attackers can disrupt the transmission and reception of semantic information.
Semantic eavesdropping attacks Physical channel Semantic information transmission Confidentiality Semantic eavesdroppers can capitalize on openness of wireless channels by employing specialized receiving equipment to capture and analyze signals in transit. If eavesdroppers obtain access to the semantic decoder, they could potentially decode some or all of the private or sensitive semantic information, even when their channel conditions are substantially inferior to those of the legitimate receivers.
Semantic inference attacks receiver and physical channel Semantic information transmission Confidentiality and Integrity By analyzing model intermediate information exposed, semantic inference attackers can infer whether specific data belongs to the training set members and even further reconstruct sensitive semantic information. This attack method directly threatens the privacy security of each user participating in distributed semantic learning.
Semantic jamming attacks physical channel Semantic information transmission Availability and Integrity They focuse on interfering with the semantic content of transmitted data, thereby degrading the consistency and quality of the data, and hindering the receiver’s accurate understanding of the intended message.